OpenStudy (anonymous):

A school choir has 80 students; 36 of them are in seventh grade. Also, 40 students in the choir are girls; 18 of them are in seventh grade. What is the probability that a student picked at random from the choir is either a girl or is in seventh grade?

OpenStudy (sleepyjess):

(girls+7th grade)/80 \((40+36)/80\) 76/80 .95 move the decimal to the right 2 places 95%

OpenStudy (anonymous):

@sleepyjess Yes I have done this but the answer choices confuse me. A.) 27/40 B.)29/40 C.)13/40 D.)33/40 None of them equal 95%

OpenStudy (sleepyjess):

hmmm then maybe there is an issue with the question wording?

OpenStudy (sleepyjess):

@iGreen @perl

OpenStudy (anonymous):

Thats what i thought too

OpenStudy (sleepyjess):

unless it is supposed to be 40+18 since there are 18 girls in 7th grade

OpenStudy (sleepyjess):

if we do 40+36 then we are counting the girls in 7th grade twice

OpenStudy (igreen):

Well..it says there are only 18 girls in 7th grade..

OpenStudy (sleepyjess):

40+16/80 56/80 28/40 70%

OpenStudy (igreen):

40 - 18 = 22. 22 + 36 = 58. 58 / 80 = 0.725 = 72.5%

OpenStudy (igreen):

So what's gonna be your answer? @tyler_ralston

OpenStudy (anonymous):

40+18/80 58/80 29/40 72.5%

OpenStudy (sleepyjess):

i see where i went wrong now i did 40+16 instead of 18

OpenStudy (igreen):

You're right @sleepyjess we counted the girls in 7th grade twice..therefore, we had to subtract them. 40 - 18 = 22 girls that aren't in 7th grade. 22 girls that aren't in 7th grade + 36 7th grade students = 58 students. 58 students / 80 total students = 72.5%

OpenStudy (igreen):

Yep, so your answer is 29/40. @tyler_ralston

OpenStudy (anonymous):

THank You @iGreen and @sleepyjess

OpenStudy (igreen):

No problem!

OpenStudy (sleepyjess):

\(\Large\cal\color{magenta}{No~problem!~Happy~to~help~\ddot\smile}\)